Please help. I put on problem on a few moments ago but here is my last one and I can show you what I I have so far. I am really not good at this but am really trying.
Define a variable, write the equation and then solve.
Three times the greatest of 3 consecutive even integers exceeds twice the least by 38. Find the integers.
So I have so far on my paper (x)+(x+2)+(x+4) to represent the 3 integers right?
Then do I represent three times the greatest by 3x on the outside of the numbers or do I include all my integers by using 3 as an exponent? This is where I am confused and get stuck.
Thank you for any help.

Answer by Earlsdon(6294)

Try this:
The three consecutive even integers are: x, x+2, and x+4
3(x+4) = 2x+38 Simplify and solve for x.
3x+12 = 2x+38 Subtract 2x from both sides.
x+12 = 38 Subtract 12 from both sides.
x = 26
The three consecutive even integers are:
26, 28, and 30
Check:
3(30) = 2(26)+38
90 = 52+38
90 = 90 OK